If a certain number is increased by 5,half of the result is three fifth of the excess of 61 over that number.Find the number.
Answer by fractalier(6550)

If I understand what you wrote, we can call the unknown, x, and write
(1/2)(x + 5) = (3/5)(61-x)
Multiply everything by ten and get
5(x+5)=6(61-x)
5x+25 = 366 - 6x
Collect like terms
11x = 341
Divide by 11 and
x = 31